OpenStudy (anonymous):

@alexrobin13 Newton's second law states that F = m x a (Force is mass times acceleration). Which example would have the GREATEST acceleration? A) a 10 kg ball thrown with a 50 Newton force B) a 1 kg ball thrown with a 0.5 Newton force C) a 50 kg student launched by a 100 Newton catapult D) a 500 kg car accelerated by 1000 Newtons from the engine

OpenStudy (alexrobin13):

A) 50/10=a, a=5 B) 0.5/1=a, a=.5 C) 100/50=a, a =2 D)1000/500=a, a =2 So A) would give the largest acceleration